[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 17, 2017

Monica Bozeman - NOAA Federal monica.bozeman at noaa.gov
Wed May 17 15:32:17 UTC 2017


GOES-16 data is available in NAWIPS, please check again. Please note, that
we will be attempting this feed switch to LDM again at 1pm today. Please
report if you notice any additional outages after this time to me or in the
shift logs please! Also, PLEASE fill out the GOES-16 feedback form!!!!
https://docs.google.com/a/noaa.gov/forms/d/e/1FAIpQLSePBDbMrGa5tWKxHZSbQnywbUAdHfhNhHMYwEVQoEV4DVwh8Q/viewform

Monica


Monica Bozeman
Meteorologist/Programmer
National Hurricane Center
301-683-1553
monica.bozeman at noaa.gov
<http://www.wdtb.noaa.gov/badges/monica.bozeman>

On Wed, May 17, 2017 at 3:26 PM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:

> Here is today's summary of NHC computer operations:
> --- Craig
>
> ----------
>
> *TAFB:*
>
> 1. NCO continues to make configuration changes to the cumulus N-AWIPS
> workstation as part of the upgrade to the RHEL6 operating system. The
> workstation is currently stable. Surface Analyst Evelyn Rivera-Acevedo will
> run it through some stress tests and report to TSB (Matt Sardi) later today.
>
> 2. To address multiple GFE ingest failures of the Atlantic/NH2 domain NWPS
> wave model output, AWIPS-II Sysadmin Chris Mello is developing a new script
> that will break the large output into separate time (snapshot) files. This
> mimics the way that most grib2 data is ingested.
>
> Chris will also work with Lead Forecaster Eric Christensen to convert all
> TAFB forecasters' procedures to D2D so the SBN data feed can eventually be
> turned off.
> *HSU:*
>
> *Other items:*
>
> 1. Switching the delivery of GOES-16 satellite imagery in AWIPS-II from
> scp (ssh secure copy) to a LDM feed has caused multiple outages. Restarting
> the LDM seems to have fixed this problem. TAFB forecasters can now view the
> imagery in D2D. Currently, no GOES-16 imagery is available in N-AWIPS. TSB
> Developer Monica Bozeman will investigate.
>
> An earlier outage of the CONUS sector on May 14th was because the
> satellite was operating in mode 4, which means 10 min full-disk imagery, no
> CONUS and no mesoscale domains.
> *NHC RFCs:*
>
> 1. Version 1.2 of the Wind Speed Probability (WSP) package, which includes
> updates to the track and error distribution coefficients (years 2012-2016
> for all basins) was implemented yesterday at 1634Z (12:37 pm EDT).
>
> 2. A Request for Change (RFC) to enable DBNet alerts on NHC's dbn1/2
> servers for the GFS ensemble spread and average output files for use by the
> reconnaissance aircraft targeting program will be implemented on Tuesday,
> May 23 by the NCO Dataflow group.
>
>
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170517/8a804081/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list